The following data are available for Sellco for the fiscal year ended on January 31, 2020: Sales 1,600 units Beginning inventory 500 units @ $ 4 Purchases, in chronological order 600 units @ $ 5 800 units @ $ 6 500 units @ $ 8 Required: a. Calculate cost of goods sold and ending inventory under the cost flow assumptions, FIFO, LIFO and weighted average (using a periodic inventory system): (Round unit cost to 2 decimal places.)
Answer #1
Unit Unit Cost Total Cost
Beginning inventory 500 4 2000
Purchase 1 600 5 3000
Purchase 2 800 6 4800
Purchase 3 500 8 4000
Total 2400 13800

Calculate cost of goods sold and ending inventory

FIFO LIFO Weighted average
Cost of goods sold (2000+3000+500*6) = 8000 (4000+4800+300*5) = 10300 13800/2400*1600 = 9200
Ending inventory 13800-8000 = 5800 13800-10300 = 3500 13800-9200 = 4600